• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kuchina

    Log entry time 16:02:14 on April17,2007

    Entry number 199384

    keyword=shift summary

    Shift workers:

    Running 90 uA on 3He.

    Run summary(run#, start-stop,N,replay?,comments)
    2572,07:58-08:30, 0.8M,Y, RC problem no beam at the end
    2573,08:34-09:03, 1.2M,Y, spot move scheduled for 9am
    9:30am Riad started scheduled parity tests, with beam 10uA
    2574,09:41-09:46, 0.4M,Y, started as soon as beam is back to 90uA, ended because MCC took the beam off.
    2575,09:53-10:20, 1.2M,Y
    2576,10:20-10:47, 1.2M,Y
    2577,10:48-11:15, 1.2M,Y
    2578,11:16-11:45, 1.2M,Y
    2579,11:46-12:14, 1.2M,
    2580,12:15-12:43, 1.2M,
    2581,12:44-13:14, 1.2M,
    2582, junk, beam is off soon after run was started
    2583,14:48-15:16, 1.2M,
    2584,15:17-15:23, Right Q1 tripped to 0
    Tried to reset, rebooted IOC nothing helped by the end of the shift.